40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4762     Entry time: Mon May 23 18:10:41 2011
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: f2p filters on PRM : not good 

During the DRMI trial I noticed that the f2p filters on PRM is not quite effective (i.e. pushing PRM in POS direction makes misalignments).

I checked the f2p filters in an easy way. I pushed POS at 0.01 Hz with an amplitude of 1000 counts and looked at the oplev error signals with / without the f2p filters.

The picture below is a time series of the POS excitation, the oplev's PITCH and YAW error signals.

You can see there still is a big coupling from POS to YAW after the f2p filters were enabled. (Its supposed to be like this)

I will redo the f2p measurement on PRM.

f2p_PRM.png

ELOG V3.1.3-